OpenStudy (anonymous):

2w^-1y^-2v^-4x2v^6yx6w^2

OpenStudy (tyteen4a03):

2w^-1 * y^-2 * v^-4 * 2v^6 * y * 6w^2 (I'll skip LaTeX because OpenStudy doesn't render negative indices correctly) Remember that a^m * a ^ n = a^(m+n). Then remember that x^-1 = 1/x.

OpenStudy (anonymous):

ok thank you